A is twice as old as B, and C is 5 years younger than B. 3 years from now, A will be 3 times as old as C. Find the ages of A,B and C
Answer by Earlsdon(6294)

A = 2B (A is twice as old as B)
C = B-5 (C is 5 years younger than B)
A+3 = 3(C+3) (3 years from now, A will be three times as old as C)
Let's put it all together now:
A+3 = 3(C+3) Simplify.
A+3 = 3C+9 Subtract 3 from both sides.
A = 3C+6 Now substitute A = 2B and C = B-5.
2B = 3(B-5)+6 Simplify.
2B = 3B-15+6
2B = 3B-9 Add 9 to both sides.
2B+9 = 3B Subtract 2B from both sides.
9 = B or B = 9
A = 2B Substitute B = 9.
A = 2(9)
A = 18
C = B-5 Substitute B = 9.
C = 9-5
C = 4
The ages are:
A is 18 years.
B is 9 years.
C is 4 years.